OpenStudy (anonymous):

For a particular angle θ, the cosine function has the following values within one cycle of the function: cos 0=3, cos π/4=0, cos π/2=-3, cos 3 π/4 = 0, cos π=3 What is the rule of the cosine function?

OpenStudy (e.mccormick):

Huh? "What is the rule of the cosine function?" The only way I know how to read that is as an alternate form of, "What is the law of cosines?" Seems odd. OK. Well, if I had to guess... based on those values they have given you specific inputs and what the outputs of those are. You probably need to define the relationship, which would be to make a function that gets those answers.

OpenStudy (dumbcow):

general cosine equation: \[y = Acos(Bx)\] from the unit circle we know that cos(0) = 1 A = 3 also we know that cos(pi) = -1 3 = 3cos(Bpi) 1 = cos(Bpi) thus B = 2 \[ y = 3\cos(2x)\]